# Mitigating the sign problem by quantum computing

### v2

@@ I INTRODUCTION
Quantum Monte Carlo (QMC) simulations are powerful and unbiased numerical tools for nonperturbative exploration of strongly correlated systems [1], which are often analytically intractable. In such approaches, the partition function of an interacting system is written as a sum (or integral) over the exponentially large configurational space in a chosen basis. Through importance sampling over a small yet representative subset of the entire configuration space, QMC algorithms can provide an efficient method to investigating large-scale many-body systems in two or more dimensions with high accuracy.
@@
In this context, Tan et al. [20] recently proposed a novel implementation of the stochastic series expansion (SSE) QMC algorithm on a quantum computer, highlighting its significant advantages over classical counterparts. By leveraging quantum superposition, this quantum computing SSE (qc-SSE) approach enables efficient evaluation of matrix elements, thereby dispensing with the no-branching requirement that typically constrains classical implementations. Moreover, the choice of basis states is no longer restricted, provided they can be efficiently prepared on a quantum computer. Building on these computational conveniences, Ref. [20] propose a general framework aimed at resolving the sign problem via quantum computing. They show that configuration weights can be rendered non-negative by adding a sufficiently large constant $M$ to each term in the Hamiltonian. Although this modification may violate the no-branching condition, it poses no obstacle to the evaluation of configuration weights within quantum algorithms. This conceptual shift could offer a practical resolution to the sign problem and broaden the applicability of the algorithm to a wider class of physical systems. To validate their algorithm, the authors simulated a one-dimensional Ising spin chain Hamiltonian with $N=3$, 4, and 5 sites. The basis states were constructed as product states of Pauli $Z$ eigenstates, transformed via non-Clifford $T$ gates, resulting in quantum configurations that are classically intractable. In all cases examined, the average energy computed using this qc-SSE approach was found to converge to the exact finite-temperature energy obtained via exact diagonalization, thereby demonstrating the efficacy of the algorithm.
@@
Although the qc-SSE approach does not eliminate the sign problem entirely, it can still serve as a generic mitigation strategy by significantly reducing the occurrence of negative configuration weights. For illustration, we consider an antiferromagnetic XY spin chain and investigate the optimal choice of the constant $M$ for alleviating the sign problem. Here we employ an operator contraction method to enhance computational efficiency (see Appendix). Without this technique, the simulations are restricted to much smaller system sizes owing to the prohibitive cost of evaluating long operator strings. Our results show that increasing $M$ significantly improves the average sign $\langle\textrm{sgn}\rangle$, which serves as a measure of the severity of the sign problem. However, this improvement comes at the cost of increased statistical errors in energy measurements, due to the increase in length of operator strings. Our findings suggest a practical trade-off, with $M=1$ offering a good balance between mitigating the sign problem and maintaining statistical accuracy. It is important to emphasize that the sign problem persists under this qc-SSE approach, particularly at low temperatures and in larger systems. Its severity also depends sensitively on system parameters, such as anisotropy. Consequently, the sign problem remains a fundamental obstacle to accurate and efficient simulation of quantum many-body systems.
The remainder of this paper is organized as follows. The conventional SSE QMC method and the sign problem are briefly reviewed in Sec. II. We outline the framework of the qc-SSE in Sec. III, where its failure in resolving the sign problem is explained. In Sec. IV, we introduce our model and then present our QMC results to show the $M$ dependence of the severity of the sign problem. The dependence of the average sign $\langle\textrm{sgn}\rangle$ on system sizes, temperature, and spin anisotropy are examined. We conclude our paper in Sec. V. The employed operator contraction method is described in Appendix.
@@ II SSE QMC and sign problem
In the qc-SSE approach, we sample the partition function by computing configuration weights using quantum circuits that represent the operator strings. This computation becomes increasingly demanding as the number of qubits grows, particularly when simulated on classical hardware. As discussed in the previous section, evaluating the configuration weight requires introducing an ancilla qubit for each term in the operator string. Consequently, the total number of qubits in the quantum circuit equals the operator string length $n$ plus the system size $N$, i.e., $N+n$ qubits are needed to implement the computation. When the operator string becomes long, the associated computational cost can become prohibitively large, especially in classical emulation scenarios. To improve computational tractability, we introduce an operator contraction method (see Appendix) that compresses the operator string length without compromising accuracy. This approach markedly boosts simulation speed and allows us to investigate systems of size up to $N=7$ that would otherwise be computationally prohibitive.
@@ IV.1 Optimizing shift constants $M_{x}$ and $M_{z}$
As in Ref. [20], our simulations employ basis states formed from product states of Pauli $Z$ eigenstates, modified by non-Clifford $T$ gate transformations. In this representation, both components of the Hamiltonian in Eq. (15) contribute off-diagonal elements, which can lead to negative configuration weights irrespective of the system size $N$. As discussed above, introducing the constants $M_{x}$ and $M_{z}$ can mitigate the sign problem, though it does not eliminate it entirely. Here, we systematically investigate the impact of these parameters on the average sign and the precision of total energy measurements in the antiferromagnetic XY spin chain.
@@
Given the observed monotonic improvement of the average sign with increasing $M$, it appears that larger values of $M$ are advantageous within the qc-SSE framework. Nonetheless, a different perspective arises when considering the absolute percentage error $|\delta E/E_{\mathrm{ED}}|$, with $\delta E$ representing the standard deviation of the simulated energy [as defined in Eq. (13)] and $E_{\mathrm{ED}}$ denoting the exact diagonalization benchmark. Fig. 1(b) reveals that for $M>1$, the absolute percentage error $|\delta E/E_{\mathrm{ED}}|$ increases with $M$. This behavior arises from $\delta E\propto\delta\langle n\rangle$ [as indicated by Eq. (13)], combined with the well-known result $\delta\langle n\rangle\propto\sqrt{\langle n\rangle}$ [21, 22]. Due to the monotonic dependence of $\langle n\rangle$ on $M$ as seen from the inset of Fig.1(a), larger values of $M$ induce greater fluctuations in energy estimates, thereby reducing accuracy despite improved average sign.
On the other hand, a small value of average sign $\langle\textrm{sgn}\rangle$ for $M<1$ also leads to large statistical errors as shown in Fig. 1(b). This trade-off is further highlighted by the energy difference $\Delta E=|E-E_{\mathrm{ED}}|$ in the insets of Fig. 1(b): at $M=1$, the estimated energy agrees well with exact diagonalization with small uncertainty, whereas for $M$ either larger than or much smaller than 1, growing statistical errors lead to noticeable discrepancies.
@@ IV.2 Dependence on system sizes and temperatures
Figure 3: (a) Size dependence of average sign $\langle\textrm{sgn}\rangle$ with fixed $T=2$; (b) temperature dependence of average sign $\langle\textrm{sgn}\rangle$ with fixed $N=3$. Here we take $M_{x}=M_{z}=1$. The inset of (a) shows the average operator string length $\langle n\rangle$.
@@
Figure 3(a) presents our data on the system-size dependence of the average sign $\langle\textrm{sgn}\rangle$ at temperature $T=2$, for system sizes up to $N=7$. Notably, results for $N>4$ were made accessible within reasonable computation time through the use of the operator contraction technique detailed in the Appendix. As seen from our findings, the average sign for $N>3$ decreases with increasing system size $N$, accompanied by a pronounced even–odd oscillatory behavior.
The decay in $\langle\textrm{sgn}\rangle$, shown separately for both even and odd $N$, aligns with the exponential suppression predicted by Eq. (5). The observed even–odd effect, wherein systems with odd $N$ exhibit smaller average signs than those with even $N$, can be attributed to geometric phases associated with the chordless cycles in the Hamiltonian [23]. This can be easily understood when working in the $z$-basis, where only the $X_{i}X_{i+1}$ terms act as off-diagonal operators and contribute negative matrix elements to the configuration weights.
To restore the original configuration $|\alpha\rangle$, certain operator strings generate cyclic paths in the configuration space, wherein the off-diagonal $X_{i}X_{i+1}$ bonds cover the chain an odd number of times. For odd $N$, these operator strings contain an odd number of $X_{i}X_{i+1}$ terms, resulting in negative contributions to the configuration weights in Eq. (3). This lead to sign cancellations and yield $\langle\textrm{sgn}\rangle<1$ for odd-length chains. In contrast, for even $N$, all contributing operator strings involve an even number of $X_{i}X_{i+1}$ terms, ensuring that the configuration weights are non-negative. Consequently, systems with even $N$ are sign-problem-free in the $z$-basis, yielding $\langle\textrm{sgn}\rangle=1$. This even-odd effect persists in the present basis choice, as clearly reflected in our numerical data. The occurrence of negative weights in even $N$ chains is less frequent, though not entirely eliminated, compared to odd $N$ chains.
@@ IV.3 Anisotropy dependence
Figure 4: Anisotropy dependence of average sign $\langle\textrm{sgn}\rangle$ for $N=3$ and $M_{z}=M_{x}=1$ at $T=2$. The inset shows the average operator string length $\langle n\rangle$.
@@
Although the qc-SSE framework can be employed to alleviate the sign problem for arbitrary Hamiltonians and in general bases, its circuit depth, as indicated by $\langle n\rangle$, is found to scale with the inverse temperature $\beta$ and system size $N$. As a result, contributions from increasingly long operator strings become significant at low temperatures and for large systems, which limits the practical applicability of the method. Consequently, it remains an open question whether the qc-SSE framework offers a substantial advantage over the classical SSE approach, given that neither method fundamentally resolves the sign problem.

### v3

@@ I INTRODUCTION
Quantum Monte Carlo (QMC) simulations are powerful and unbiased numerical tools for nonperturbative exploration of strongly correlated systems [1], which are often analytically intractable. In such approaches, the partition function of an interacting system is written as a sum (or integral) over the exponentially large configurational space in a chosen basis. Through importance sampling over a small yet representative subset of the entire configuration space, QMC algorithms can provide an efficient method to investigate large-scale many-body systems in two or more dimensions with high accuracy.
@@
In this context, Tan et al. [20] recently proposed a novel implementation of the stochastic series expansion (SSE) QMC algorithm on a quantum computer, highlighting its significant advantages over classical counterparts. By leveraging quantum superposition, this quantum computing SSE (qc-SSE) approach enables efficient evaluation of matrix elements, thereby dispensing with the no-branching requirement that typically constrains classical implementations. Moreover, the choice of basis states is no longer restricted, provided they can be efficiently prepared on a quantum computer. Building on these computational conveniences, Ref. [20] proposes a general framework aimed at resolving the sign problem via quantum computing. They show that configuration weights can be rendered non-negative by adding a sufficiently large constant $M$ to each term in the Hamiltonian. Although this modification may violate the no-branching condition, it poses no obstacle to the evaluation of configuration weights within quantum algorithms. This conceptual shift could offer a practical resolution to the sign problem and broaden the applicability of the algorithm to a wider class of physical systems. To validate their algorithm, the authors simulated a one-dimensional Ising spin chain Hamiltonian with $N=3$, 4, and 5 sites. The basis states were constructed as product states of Pauli $Z$ eigenstates, transformed via non-Clifford $T$ gates, resulting in quantum configurations that are classically intractable. In all cases examined, the average energy computed using this qc-SSE approach was found to converge to the exact finite-temperature energy obtained via exact diagonalization, thereby demonstrating the efficacy of the algorithm.
@@
Although the qc-SSE approach does not eliminate the sign problem entirely, it can still serve as a generic mitigation strategy by significantly reducing the occurrence of negative configuration weights. For illustration, we consider an antiferromagnetic XY spin chain and investigate the optimal choice of the constant $M$ for alleviating the sign problem. Here we employ an operator contraction method to enhance computational efficiency (see Appendix A). Without this technique, the simulations are restricted to much smaller system sizes owing to the prohibitive cost of evaluating long operator strings. Our results show that increasing $M$ significantly improves the average sign $\langle\textrm{sgn}\rangle$, which serves as a measure of the severity of the sign problem. However, this improvement comes at the cost of increased statistical errors in energy measurements, due to the increase in length of operator strings. Our findings suggest a practical trade-off, with $M=1$ offering a good balance between mitigating the sign problem and maintaining statistical accuracy. It is important to emphasize that the sign problem persists under this qc-SSE approach, particularly at low temperatures and in larger systems. Its severity also depends sensitively on system parameters, such as anisotropy. Consequently, the sign problem remains a fundamental obstacle to accurate and efficient simulation of quantum many-body systems.
The remainder of this paper is organized as follows. The conventional SSE QMC method and the sign problem are briefly reviewed in Sec. II. We outline the framework of the qc-SSE in Sec. III, where its failure in resolving the sign problem is explained. In Sec. IV, we introduce our model and then present our QMC results to show the $M$ dependence of the severity of the sign problem. The dependence of the average sign $\langle\textrm{sgn}\rangle$ on system sizes, temperature, and spin anisotropy are examined. We conclude our paper in Sec. V. The employed operator contraction method is described in Appendix A.
@@ II SSE QMC and sign problem
In the qc-SSE approach, we sample the partition function by computing configuration weights using quantum circuits that represent the operator strings. This computation becomes increasingly demanding as the number of qubits grows, particularly when simulated on classical hardware. As discussed in the previous section, evaluating the configuration weight requires introducing an ancilla qubit for each term in the operator string. Consequently, the total number of qubits in the quantum circuit equals the operator string length $n$ plus the system size $N$, i.e., $N+n$ qubits are needed to implement the computation. When the operator string becomes long, the associated computational cost can become prohibitively large, especially in classical emulation scenarios. To improve computational tractability, we introduce an operator contraction method (see Appendix A) that compresses the operator string length without compromising accuracy. This approach markedly boosts simulation speed and allows us to investigate systems of size up to $N=7$ that would otherwise be computationally prohibitive.
@@ IV.1 Optimizing shift constants $M_{x}$ and $M_{z}$
As in Ref. [20], our simulations employ basis states formed from product states of Pauli $Z$ eigenstates, modified by a Hadamard gate and non-Clifford $T$ gate transformations. In this representation, both components of the Hamiltonian in Eq. (15) contribute off-diagonal elements, which can lead to negative configuration weights irrespective of the system size $N$. As discussed above, introducing the constants $M_{x}$ and $M_{z}$ can mitigate the sign problem, though it does not eliminate it entirely. Here, we systematically investigate the impact of these parameters on the average sign and the precision of total energy measurements in the antiferromagnetic XY spin chain.
@@
Given the observed monotonic improvement of the average sign with increasing $M$, it appears that larger values of $M$ are advantageous within the qc-SSE framework. Nonetheless, a different perspective arises when considering the absolute percentage error $|\delta E/E_{\mathrm{ED}}|$, with $\delta E$ representing the standard deviation of the simulated energy [as defined in Eq. (13)] and $E_{\mathrm{ED}}$ denoting the exact diagonalization benchmark. Fig. 1(b) reveals that for $M>1$, the absolute percentage error $|\delta E/E_{\mathrm{ED}}|$ increases with $M$. This behavior arises from $\delta E\propto\delta\langle n\rangle$ [as indicated by Eq. (13)], combined with the well-known result $\delta\langle n\rangle\propto\sqrt{\langle n\rangle}$ [21, 22]. Due to the monotonic dependence of $\langle n\rangle$ on $M$ as seen from the inset of Fig.1(a), larger values of $M$ induce greater fluctuations in energy estimates, thereby reducing accuracy despite improved average sign. On the other hand, a small value of average sign $\langle\textrm{sgn}\rangle$ for $M<1$ also leads to large statistical errors as shown in Fig. 1(b). This trade-off is further highlighted by the energy difference $\Delta E=|E-E_{\mathrm{ED}}|$ in the insets of Fig. 1(b): at $M=1$, the estimated energy agrees well with exact diagonalization with small uncertainty, whereas for $M$ either larger than or much smaller than 1, growing statistical errors lead to noticeable discrepancies.
@@ IV.2 Dependence on system sizes and temperatures
Figure 3: (a) Size dependence of average sign $\langle\textrm{sgn}\rangle$ with fixed $T=2$; (b) temperature dependence of average sign $\langle\textrm{sgn}\rangle$ with fixed $N=3$. Here we take $M_{x}=M_{z}=1$. The inset of (a) shows the average operator string length $\langle n\rangle_{|W|}$.
@@
Figure 3(a) presents our data on the system-size dependence of the average sign $\langle\textrm{sgn}\rangle$ at temperature $T=2$, for system sizes up to $N=7$. Notably, results for $N>4$ were made accessible within reasonable computation time through the use of the operator contraction technique detailed in the Appendix A. As seen from our findings, the average sign for $N>3$ decreases with increasing system size $N$, accompanied by a pronounced even–odd oscillatory behavior.
The decay in $\langle\textrm{sgn}\rangle$, shown separately for both even and odd $N$, aligns with the exponential suppression predicted by Eq. (5). The observed even–odd effect, wherein systems with odd $N$ exhibit smaller average signs than those with even $N$, can be attributed to geometric phases associated with the chordless cycles in the Hamiltonian [23]. This can be easily understood when working in the $z$-basis, where only the $X_{i}X_{i+1}$ terms act as off-diagonal operators and contribute negative matrix elements to the configuration weights. To restore the original configuration $|\alpha\rangle$, certain operator strings generate cyclic paths in the configuration space, wherein the off-diagonal $X_{i}X_{i+1}$ bonds cover the chain an odd number of times. For odd $N$, these operator strings contain an odd number of $X_{i}X_{i+1}$ terms, resulting in negative contributions to the configuration weights in Eq. (3). This leads to sign cancellations and yield $\langle\textrm{sgn}\rangle<1$ for odd-length chains. In contrast, for even $N$, all contributing operator strings involve an even number of $X_{i}X_{i+1}$ terms, ensuring that the configuration weights are non-negative. Consequently, systems with even $N$ are sign-problem-free in the $z$-basis, yielding $\langle\textrm{sgn}\rangle=1$. This even-odd effect persists in the present basis choice, as clearly reflected in our numerical data. The occurrence of negative weights in even $N$ chains is less frequent, though not entirely eliminated, compared to odd $N$ chains.
@@ IV.3 Anisotropy dependence
Figure 4: Anisotropy dependence of average sign $\langle\textrm{sgn}\rangle$ for $N=3$ and $M_{z}=M_{x}=1$ at $T=2$. The inset shows the average operator string length.
@@
We note that the present approach alleviates the sign problem more effectively than the standard classical SSE, as illustrated in Appendix B. In addition, the classical framework offers no general strategy for eliminating negative weights across arbitrary Hamiltonians. A traditional approach is to search for a basis in which the Hamiltonian remains non-branching and, ideally, exhibits a reduced sign problem [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. However, changing the basis typically requires rewriting the directed-loop equations and modifying the loop-update algorithm, which becomes increasingly complicated when multiple spins are coupled in the new basis. In contrast, qc-SSE entails no substantial algorithmic modifications and can be applied directly to general Hamiltonians in any basis, thereby mitigating negative signs regardless of branching.
Although the qc-SSE framework can be employed to alleviate the sign problem for arbitrary Hamiltonians and in general bases, its circuit depth, as indicated by $\langle n\rangle$, is found to scale with the inverse temperature $\beta$ and system size $N$. As a result, contributions from increasingly long operator strings become significant at low temperatures and for large systems, which limits the practical applicability of the method.
A promising direction for further improvement lies in optimizing the choice of basis states used to construct the quantum circuits. Since the severity of the sign problem is inherently basis dependent, variationally tuning the local basis—such as optimizing the orientation of single-spin or multi-spin basis states—may further suppress negative-weight contributions without introducing additional algorithmic complexity. Within the qc-SSE framework, such basis optimization can be implemented at the circuit level without modifying the Monte Carlo update structure.
@@ Appendix A Operator contraction
Appendix B Comparison with classical SSE
Figure 5: Average sign $\langle\textrm{sgn}\rangle$ of classical SSE on z-basis and qc-SSE as a function of $T$ for $\Delta=1$ and $N=3$.
To demonstrate the advantage of using qc-SSE over the standard classical SSE, we compare the average sign $\langle\textrm{sgn}\rangle$ obtained from both methods as a function of temperature $T$ for the case of $\Delta=1$ and $N=3$. For qc-SSE, we set $M_{z}=M_{x}=1$. By contrast, for the classical SSE on $z$-basis, $M_{z}=1$ but $M_{x}$ must be set to zero, since any finite $M_{x}$ introduces branching processes that the classical SSE cannot handle. As shown in Fig. 5, the average sign produced by the proposed qc-SSE remains significantly higher than that of the classical SSE, highlighting the improved sign performance of the qc-SSE approach. It can be understood that introducing a finite $M_{z}$ and $M_{x}$ suppresses the negative contribution arising from the $Z_{i}Z_{i+1}$ and $X_{i}X_{i+1}$ operators, thereby allievating the sign problem.